OpenStudy (anonymous):

sum(5n^n/n!) as n -> infinity

OpenStudy (anonymous):

not much chance of that the terms get larger, not smaller

OpenStudy (anonymous):

how do you prove it using the limit test?

OpenStudy (anonymous):

Geometric sequences: Determine whether the following sequences converge or diverge & describe whether they do so monotonically or by oscillation. {0.2^n}

OpenStudy (anonymous):

limit test just says if \(\lim_{n\to \infty}a_n\neq 0\) then the sum does not converge

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

is it clear that the terms do not go to zero?

OpenStudy (anonymous):

if not try it with say \(n=5\) and see what you have exactly \[\frac{5\times 5\times 5\times 5\times 5}{5\times 4\times 3\times 2\times 1}\]\[=\frac{5}{4}\times \frac{5}{3}\times \frac{5}{2}\times 5\] so in fact this grow rather fast, faster than \(n\)

OpenStudy (anonymous):

right okay i was a little confused on proving that lim n→∞ an≠0